# Staking out the Proton Drip-Line of Thulium at the N=82 Shell Closure

### v1

@@ Abstract
Direct observation of proton emission with very small emission energy is often unfeasible due to the long partial half-lives associated with tunneling through the Coulomb barrier. Therefore proton emitters with very small Q-values may require masses of both parent and daughter nuclei to establish them as proton unbound. Nuclear mass models have been used to predict the proton drip-line of the thulium (Tm) isotopic chain ($Z=69$), but up until now the proton separation energy has not been experimentally tested. Mass measurements were therefore performed using a Multiple Reflection Time-Of-Flight Mass Spectrometer [1] (MR-TOF-MS) at TRIUMF’s TITAN facility to definitively map the limit of proton-bound Tm. The masses of neutron-deficient, 149Tm and 150Tm, combined with measurements of 149m,gEr (which were found to deviate from literature by $\sim$150 keV), provide the first experimental confirmation that 149Tm is the first proton-unbound nuclide in the Tm chain. Our measurements also enable the strength of the $N=82$ neutron shell gap to be determined at the Tm proton drip-line, providing evidence supporting its continued existence.
@@ I Introduction
In order for a proton to be emitted, it first must tunnel through the Coulomb barrier (i.e. a single-nucleon analogue to $\alpha$-decay). In contrast to alpha decay, it is however simpler to calculate the proton emission decay rate or spectroscopic factors as it doesn’t require the formation of a cluster of nucleons [2]. On the other hand, proton emission is significantly influenced by the centrifugal barrier and the proton emission half-life at a given proton number, $Z$, depends strongly on the underlying nuclear structure, the angular momentum carried away by the emitted proton [2][4][5], and the decay energy, $Q_{p}$ [5].
Furthermore, it has become evident that the nuclear shell structure can change towards the drip-line. Effects such as the weakening or disappearance of shells (quenching) at the classical magic numbers and the appearance of new magic numbers have been predicted and observed [6, 7]. Nuclear masses and binding energies provide indispensable information when studying the limits of nuclear existence. They allow calculations of the particle separation energies and can be used to make predictions regarding the underlying nuclear structure and nuclear properties such as half-life, angular momentum, or spectroscopic factors in the case of proton emission. Moreover, the proton drip-line around $Z=69$ is associated with an abrupt change in the deformation [8, 2, 9].
The mass values of nuclides across the $N=82$ shell closure, formed between the $h_{11/2}$ (or nearby $s_{1/2}$ or $d_{3/2}$) and the $f_{7/2}$ orbitals, have been intensively studied on the neutron-rich side of the nuclear chart, see e.g. [10, 11, 12, 13, 14, 15]. Corresponding investigations at the neutron-deficient $N=82$ shell closure up to the elements of holmium ($Z=67$) and erbium ($Z=68$) have been preformed by SHIPTRAP [16], ISOLTRAP [17], by the ESR using Schottky Mass Spectrometry [18] and recently in ytterbium ($Z=70$) at TITAN [19]. However masses of neutron-deficient thulium ($Z=69$) isotopes across $N=82$ remain elusive right where the one-proton drip-line is predicted to lie.
Here we report first time high precision mass measurements of neutron-deficient thulium isotopes, which close the gap across the proton drip-line and $N=82$ in the isotopic chain.
@@ I.1 Background
Ground-state proton emission can occur for any nucleus beyond the proton drip-line, and was first observed in 1982 in 151Lu and 147Tm [20, 21]. Since then, at least 27 heavy ground-state proton emitters have been identified, alongside a number of isomeric states [8].
While 147Tm has long been known to be a proton emitter, and prior direct mass measurements of 147Tm and 148Tm have been able to constrain the location of the proton drip line to lie somewhere between $A=148$ and $A=151$ [16], the precise isotope at which Tm becomes proton-unbound remains uncertain due to the thus far unmeasured mass values of 149Tm and 150Tm. In cases where neither observation of the decay nor direct mass measurements can be used to establish the drip-line, various theoretical approaches are employed to predict its location but these models are known to struggle to predict $S_{p}$ far from stability [22]. Density functional theory and phenomenological descriptions of the nuclear binding energy have been utilized to predict the location of the proton drip-line. For instance, the Skyrme results of [23] can be used to interpolate the $S_{p}$ of odd-Z nuclei. However, for 149Tm, the choice of energy density functional – such as UNEDF1[24] or SV-min[25] – leads to competing predictions regarding its position relative to the drip-line. Furthermore, the newer BSkG3 model [26] places the drip-line even closer to stability.
Bayesian model averaging can be used to assign weights to models in an attempt to improve estimates of the binding energies of heavy nuclei [27]. A recent Bayesian statistical analysis was performed to weight the proton drip-line predictions provided by several different energy density functionals, as well as FRDM2012 and HFB-24 [28]. This work predicted the last proton-bound isotope of Tm to occur at $N=81$ ($A=150$), but this has never been experimentally verified.The precise location of the Tm drip-line therefore remains a significant hole in our experimental knowledge of the heavy, proton-deficient nuclei, and the present measurement aims to uncover the exact location of the ground-state proton drip-line of Tm ($Z=69$).
@@ II Experimental Description
The isotopes under study were produced throughout two experimental campaigns at TRIUMF’s Isotope Separator and Accelerator (ISAC) facility[29] using the ISOL technique. A 480 MeV proton beam was impinged on a Ta target to produce rare isotopes via spallation reactions and Tm and Er isotopes were surface ionized in the target region. The continuous beam of neutron-deficient isotopes was purified using the ISAC high-resolution mass separator [30] (resolving power $\sim$2500) to select beams from $A=149$ to $A=157$ before preparation in the TITAN Radio-Frequency Quadrupole (RFQ) cooler-buncher trap[31] for precision mass measurements. The resulting bunches of ions were extracted and their kinetic energy matched for injection into TITAN’s MR-TOF-MS [32][33][34] to obtain a mass spectrum. In the MR-TOF-MS ions are again cooled and narrowly bunched before executing isochronous (energy independent) reflections between two electrostatic ion mirrors and being detected using a MagneTOFTM ion detector manufactured by ETP ion detect. All ions observed in the spectrum were delivered and measured in a 1+ charge state and any possible non-isobars present due to charge exchange were removed using a Mass Range Selector [35, 36]. An example spectrum of $A=149$ is given in Figure 1.
Measurements of the nuclei furthest from stability were made possible by employing so called mass selective re-trapping [37, 38] to selectively reduce the intensity of unwanted contaminant species [19]. This improved the signal-to-background ratio of the 149,150Tm measurements and allowed the MR-TOF-MS to accept a higher overall beam rate while keeping the total number of ions detected per mass measurement cycle below one.
@@
The mass of 150Tm was measured alongside Yb masses [19] in experiment #1 which additionally provided anchor masses for alpha decay chains [39]. The resolving power of 270k in this experiment resulted in a proton separation energy of 150Tm which was very close to zero, and no definitive conclusion could be drawn as to whether this nucleus is proton-bound. The Tm data-set was greatly improved during experiment #2 by performing additional measurements extending measurements to 149Tm, and improving on the initial experimental uncertainties. Particularly, experiment #2 benefited from improvements that were made to the MR-TOF-MS system to achieve 400k resolving power and lower systematic uncertainty.
Figure 1: Top: The $A=149$ spectrum with major peaks labeled. 149Dy is used as both shape and mass calibrant here ($m$ indicates a long-lived isomer). Bottom: A fit of the 149g+mEr peak using the hG functional form and calibrated to the shape of the higher statistics 149Dy peak. Use of the hyperEMG model enabled the mass of the ground-state to be directly determined alongside the isomer within the mixed peak.
@@ II.1 Analysis Procedure
The analysis procedure followed the description in [40]. At all mass units, we utilized a high abundance peak to correct for time-dependent drifts in the Time of Flight (TOF) spectrum using in-house developed data acquisition software[33]. Subsequently, masses were determined by fitting the peaks in each spectrum using a hyper-exponentially-modified Gaussian (hyperEMG) lineshape[41]. This procedure involves first fitting an isolated high-statistics peak with a multi-dimensional function to precisely determine the shape (i.e. functional form) of a single peak (the shape calibrant) and then using multiple copies of this shape to fit all peaks in the spectrum. Next, the masses of all peaks in the spectrum were calibrated against a single species with a well-established mass (the mass calibrant). Appropriate shape and mass calibrant peaks were chosen for each set of isobars measured to minimize the influence of overlapping isobars and nuclear isomers. In the present analysis, the peaks corresponding to 149Dy, 150Ho, 151Ho, 152Dy, 153Dy, 154Dy, 155Yb, 156Yb, and 157Yb were selected as shape calibrants in their respective isobaric mass spectra. The chosen mass calibrant at each mass number was cross-checked for consistency with the masses of several different isobars listed in the 2020 Atomic Mass Evaluation (AME2020) to confirm its identity, and are given in Table 1. However, multiple species contained long lived isomers which were not resolvable from their respective ground-state. Such unresolved contributions may shift the resulting mass values to higher mass as would be expected for a pure ground-state of the given isotope. When the excitation energy is known, we apply a correction and additional uncertainty following the guidelines laid out in the AME2020 appendix for the treatment of unresolved peaks. However, according to ENSDF ${}^{151,152,154}_{~{}82,~{}83,~{}85}$m respectively contain a $\sim 7$ s, a $5.2$ s, and a $3.3$ s isomer of unknown excitation energy which are likely to have influenced the measurement [42, 43, 44]. The mass values corresponding to the mixed peaks are included for completeness.
@@ II.2 Results
The final mass values were obtained by adding statistical, fit and known systematic uncertainties in quadrature and are summarized in Table 1. Mass excess (ME) uncertainties include the uncertainty of the calibrant mass, while the Ion Mass Ratio uncertainties do not. The systematic uncertainty was dominated by non-ideal switching of voltages in the second ion mirror. For the initial measurement of 150Tm presented here a systematic uncertainty of $3\times 10^{-7}$ was included, whereas for all other measurements obtained during experiment #2 a systematic uncertainty of $1.5\times 10^{-7}$ was incorporated. This reduction was the result of improvements made to the tune of the MR-TOF-MS and the stability of the high voltage electrodes.
@@
The crowded $A=153$ spectrum was the most challenging from which to obtain a high-statistics shape calibrant in the analysis. An isobaric species lies in each of the two tails of the 153Dy shape/mass calibrant, and the functional form needed to be manually selected in order to ignore excessive influence to the fit from the isobar in the lower mass tail. Furthermore, having a total rate of $\sim$3.8 ions per trapping cycle, with $\sim$2.7 per cycle in the 153Dy calibrant peak, necessitated the inclusion of an additional systematic uncertainty of $8\times 10^{-8}$ per ion per cycle to account for ion-ion interactions.
The species 154Tm, 152Tm, and 151Tm each contain a long-lived isomer for which the excitation energy is unmeasured. Since both the excitation energy and the ratio of ground-state to isomer entering the MR-TOF-MS are unknown, it is difficult to assign a definite mass for either ground-state or isomer, but the results are included for completeness.
153Tm has a long-lived isomer at 43.2$\pm$0.2 keV and 155Tm one at 41$\pm$6 keV. Both mass peaks are likely an unresolved admixture of isomer and ground-state. As per the standard procedure in Appendix B of AME 2020[45], the ground-state mass ($m_{gs}$) was obtained from the experimental mass ($m_{exp}$) and the known excitation energy ($\Delta E$) through the relation $m_{\textrm{ground-state}}=m_{\textrm{experiment}}-\Delta E/2$, and a corresponding additional uncertainty ($\delta m_{\textrm{admixture}}$) was included, where $\delta m_{\textrm{admixture}}=0.290(\Delta E)$.
The mass of 157Lu was likewise corrected for an unknown contribution from a 20.9$\pm$2.0 keV isomer (4.8 s) to obtain a mass 56$\pm$29 keV heavier than the AME value.
The mass spectra of 149,150Tm strongly benefited from mass selective retrapping improving the Tm-to-contaminant ratio. Once applied the peaks were well separated and could be fitted. No indication of the 670 keV, 5 ms 150Tm isomer was seen in the mass spectrum.
@@
Overall we see good agreement between our new mass values and literature values from AME2020, particularly for 157Tm and 156Tm which are not known to have long-lived isomers or isobaric species that could impact our measurements. The two new Tm mass values now close the gap between a series of previously measured neutron deficient isotopes [16, 46, 47] and the neutron-rich side of the nuclear chart, with mass data now spanning from N=78 (147Tm) all the way to N=99 (168Tm). Alongside the measured Tm isotopes, 149-154Er were also identified in the mass spectra. The masses of 150-154Er all lie within 1.5 $\sigma$ of the presently accepted values in the AME2020 and any known isomers are $>2.5$ MeV above the ground state and thus do not interfere with the measurements.
@@
149gEr and 149mEr were fit using two hyperEMG functions calibrated to the 149Dy peak (Fig. 1) to account for its known $11/2^{-}$ isomeric state, which was dominantly delivered to TITAN. Our 149gEr and 149mEr masses were found to be heavier by 2.9 $\sigma$ and 3.8 $\sigma$ respectively when compared to the literature value obtained via a measurement of 149mEr at the GSI ESR [18]. This results in a shift of the $S_{p}$ of 150Tm (see below).
@@
Table 1: Mass measurements with their respective calibrants, ion mass ratios, and mass excess results compared to AME2020 values. The 150Tm mass from experiment #1 [19] is included at the bottom. The # indicates an extrapolated mass in the AME2020. Nuclides with a * have a long-lived isomer of unknown excitation energy, which may influence the measurement to an unknown degree. No correction was made and no additional uncertainty added and the mass excess of the admixture is reported (in italics). Nuclides with a contained an unresolved admixture of the ground-state and an isomer of a known excitation energy in the peak; for these a correction was made to the mass excess and the ion mass ratio (in bold) and an additional uncertainty added to reflect this (see section III). The 155Tm peak has an unresolved 41(6)keV isomer [48], 153Tm has a 43.2$\pm$0.2 keV isomer [49], and 157Lu has a 20.9$\pm$2.0 keV isomer [50]. As described in the text, these values were corrected using the procedure outlined in the 2020 Atomic Mass Evaluation (AME2020) [45] to obtain a mass for the ground-state and its uncertainty. A corresponding additional uncertainty of 13 keV for 153gTm and 12 keV for 155gTm is included. The high rate of 153Dy also necessitated the inclusion of an additional systematic uncertainty of $3\times 10^{-7}$.

### v3

# Convergence on the Proton Drip-Line in Thulium

@@ Abstract
Direct observation of proton emission for very small Q-values is often unfeasible due to the long partial half-lives of the proton emission channel associated with tunneling through the Coulomb barrier. Therefore, proton emitters with very small decay energies may require the masses of both parent and daughter nuclei in order to establish them as proton unbound. Nuclear mass models have been used to predict the proton drip-line of the thulium (Tm) isotopic chain ($Z=69$), but until now the proton separation energy has not been experimentally tested. Mass measurements were performed using a Multiple Reflection Time-Of-Flight Mass Spectrometer (MR-TOF-MS) at TRIUMF’s TITAN facility to conclusively map the limit of proton-bound Tm. The masses of neutron-deficient, 149Tm and 150Tm, combined with measurements of 149m,gEr (which were found to deviate from literature by $\sim$150 keV), provide the first experimental confirmation that 149Tm is the first proton-unbound nuclide in the Tm chain. Our measurements also enable the strength of the $N=82$ neutron shell gap to be determined at the Tm proton drip-line, providing evidence supporting its continued existence.
@@ I Introduction
Near the proton drip-line, a proton must tunnel through the Coulomb barrier in order to be emitted (i.e. a single-nucleon analogue to $\alpha$-decay). In contrast to alpha decay however, it is simpler to calculate the proton emission decay rate or spectroscopic factor, since the decay does not require the formation of a cluster of nucleons [1]. Proton emission is significantly influenced by the centrifugal barrier and the proton emission half-life at a given proton number, $Z$, depends strongly on underlying nuclear structure including the angular momentum carried away by the emitted proton [1][3][4], and the decay energy, $Q_{p}$ [4].
Furthermore, the nuclear shell structure can change towards the drip-line; effects such as the weakening or disappearance of shells (quenching) at the classical magic numbers, and the appearance of new magic numbers have been both predicted and observed [5, 6]. Nuclear masses and binding energies provide indispensable information when studying the limits of nuclear existence and shell structure. They allow calculations of the particle separation energies and can be used to make predictions regarding the underlying nuclear structure and nuclear properties such as half-life, angular momentum, or spectroscopic factors in the case of proton emission.
The mass values of nuclides across the $N=82$ shell closure, formed between the $h_{11/2}$ (or nearby $s_{1/2}$ or $d_{3/2}$) and the $f_{7/2}$ orbitals, have been intensively studied on the neutron-rich side of the nuclear chart, see e.g. [7, 8, 9, 10, 11, 12]. Corresponding investigations at the neutron-deficient $N=82$ shell closure up to the elements of holmium ($Z=67$) and erbium ($Z=68$) have been performed by ISOLTRAP [13], the ESR using Schottky Mass Spectrometry [14], and recently in ytterbium ($Z=70$) at TITAN [15]. The proton drip-line around $Z=69$ is associated with an abrupt change in the deformation [16, 1, 17]. Although masses of neutron-deficient thulium isotopes have been measured by SHIPTRAP [18], measurements remain elusive across the $N=82$ shell closure where the one-proton drip-line is predicted to lie. The MR-TOF-MS technique [19] is well suited for measurements of short-lived isotopes. Here we report the first high-precision mass measurements of neutron-deficient thulium isotopes at the proton drip-line, enabling S2N to be determined across $N=82$.
@@ I.1 Background
Ground state proton emission can occur for any nucleus beyond the proton drip-line, and was first observed in 1982 in 151Lu and 147Tm [20, 21]. Since then, at least 27 heavy ground state proton emitters have been identified, along with a number of isomeric states [22][23].
While 147Tm has long been known to be a proton emitter, and prior direct mass measurements of 147Tm and 148Tm have been able to constrain the location of the proton drip-line to lie somewhere between $A=148$ and $A=151$ [18], the precise isotope at which Tm becomes proton-unbound remains uncertain due to the thus far unmeasured masses of 149Tm and 150Tm. In cases where neither observation of the decay nor direct mass measurements can be used to establish the drip-line, various theoretical approaches are employed to predict its location, but these models are known to struggle to predict $S_{p}$ far from stability [24].
Density functional theory and phenomenological descriptions of the nuclear binding energy have been utilized to predict the location of the proton drip-line. For instance, the Skyrme results of [25] can be used to interpolate the $S_{p}$ of odd-Z nuclei. However, for 149Tm, the choice of energy density functional – such as UNEDF1 [26] or SV-min [27] – leads to competing predictions regarding its position relative to the drip-line. Furthermore, the newer BSkG3 model [28] predicts the drip-line to be even closer to stability.
Bayesian model averaging can be used to assign weights to models in an attempt to improve estimates of the binding energies of heavy nuclei [29]. A recent Bayesian statistical analysis was performed to weight the proton drip-line predictions provided by several different energy density functionals, as well as FRDM2012 and HFB-24 [30]. This work predicted the last proton-bound isotope of Tm to occur at $N=81$ ($A=150$), but this has never been experimentally verified.The precise location of the thulium proton drip-line remains a significant gap in our experimental knowledge of the heavy, proton-deficient nuclei. The present measurement aims to determine the exact position of the ground state proton drip-line for Tm ($Z=69$).
@@ II Experimental Description
The isotopes under study were produced throughout two experimental campaigns at TRIUMF’s Isotope Separator and Accelerator (ISAC) facility [31] using the ISOL (Isotope Separation On-Line) technique. A 480 MeV proton beam was impinged on a Ta target to produce rare isotopes via spallation reactions and the produced Tm and Er isotopes were surface ionized. The continuous beam of neutron-deficient isotopes was purified using the ISAC high-resolution mass separator [32] (resolving power $\sim$2500) to select beams from $A=149$ to $A=157$ before preparation in the TITAN Radio-Frequency Quadrupole (RFQ) cooler-buncher trap [33] for precision mass measurements. The resulting bunches of ions were extracted and their kinetic energy matched for injection into TITAN’s MR-TOF-MS [34][35][36] to obtain a mass spectrum. Ions were again cooled and narrowly bunched in the injection trap of the MR-TOF-MS, before they underwent isochronous (energy independent) reflections between two electrostatic ion mirrors before time-of-flight detection using a MagneTOFTM ion detector manufactured by ETP ion detect. All ions observed in the spectrum were delivered and measured in a 1+ charge state and any possible non-isobars present due to charge exchange were removed using a Mass Range Selector [37, 38]. An example spectrum of $A=149$ is given in Figure 1.
Measurements of the nuclei furthest from stability were made possible by employing so-called mass selective re-trapping [39, 40], during which the ion bunch is captured a second time in the injection trap after a first stage of mass separation in the MR-TOF-MS. Mass selective re-trapping can be used to selectively reduce the intensity of the dominant contaminant species by more than three orders of magnitude [15]. This improved the signal-to-background ratio of the 149,150Tm measurements and allowed the MR-TOF-MS to accept a higher overall beam rate while keeping the total number of ions detected per mass measurement cycle within an acceptable range.
@@
The mass of 150Tm was measured along with the Yb masses [15] obtained in experiment #1, which additionally provided anchor masses for alpha decay chains [41]. The resolving power of $\sim$270,000 in this experiment resulted in a proton separation energy of 150Tm which was very close to zero, and no definitive conclusion could be drawn as to whether this nucleus is proton-bound. The Tm data-set was greatly improved during experiment #2 by performing additional measurements which extended data to 149Tm, and improved on the initial experimental uncertainties. In particular, experiment #2 benefited from improvements that were made to the MR-TOF-MS system to achieve $\sim$400,000 resolving power and lower systematic uncertainty.
Figure 1: Top: The $A=149$ spectrum with major peaks labeled. 149Dy is used as both shape and mass calibrant here ($m$ indicates a long-lived isomer). Bottom: A fit of the 149g+mEr peak using the hyperEMG functional form and calibrated to the shape of the higher statistics 149Dy peak. The known peak shape enabled the mass of the ground state to be directly determined alongside the isomer within the mixed peak.
@@ II.1 Analysis Procedure
The analysis procedure followed the description in [42]. At all mass units, we utilized a high abundance peak to correct for time-dependent drifts in the Time of Flight (TOF) spectrum using in-house developed data acquisition software [35]. Subsequently, masses were determined by fitting the peaks in each spectrum using a hyper-exponentially-modified Gaussian (hyperEMG) lineshape [43] and the emgfit python library available at [44]. This procedure involves first fitting an isolated high-statistics peak with a multi-dimensional function to precisely determine the shape (i.e. functional form) of a single peak (the shape calibrant) and then using multiple copies of this shape to fit all peaks in the spectrum. Next, the masses of all peaks in the spectrum were calibrated against a single species with a well-established mass (the mass calibrant). Appropriate shape and mass calibrant peaks were chosen for each set of isobars measured to minimize the influence of overlapping isobars and nuclear isomers. In the present analysis, the peaks corresponding to 149Dy, 150Ho, 151Ho, 152Dy, 153Dy, 154Dy, 155Yb, 156Yb, and 157Yb were selected as shape calibrants in their respective isobaric mass spectra. The chosen mass calibrant at each mass number was cross-checked for consistency with the masses of several different isobars listed in the 2020 Atomic Mass Evaluation (AME2020) to confirm its identity, and are listed in Table 1.
@@ II.2 Results
The final mass values were obtained by adding the experimental and known systematic uncertainties in quadrature and making corrections for any known or expected unresolved isomers. These are summarized in Table 1. Additionally, Table 2 contains the uncorrected measurement data – excluding any isomer corrections – and is provided as a reference should new information about these isomers become available in the future. Mass excess (ME) uncertainties include the uncertainty of the calibrant mass, while the Ion Mass Ratio uncertainties do not. The systematic uncertainty was dominated by nonideal switching of voltages in the ion mirror on the extraction side, and was measured using stable ions as described in [45] and [42]. The single measurement presented here from experiment #1, 150Tm, includes a systematic uncertainty of $3\times 10^{-7}$, while the measurements obtained during experiment #2 incorporate a systematic uncertainty of $1.5\times 10^{-7}$ (except A=153). This reduction was the result of improvements made to the tune of the MR-TOF-MS and the stability of the high-voltage electrodes.
@@
The crowded $A=153$ spectrum shown in Figure 2 was the most challenging from which to obtain a high-statistics shape calibrant in the analysis. An isobaric species lies in each of the two tails of the 153Dy shape/mass calibrant, and the functional form needed to be manually selected in order to ignore excessive influence on the fitting algorithm from the isobar in the lower mass tail. Furthermore, the mass 153 spectrum had a total rate of $\sim$3.8 ions per trapping cycle, with $\sim$2.7 per cycle in the 153Dy calibrant peak. Therefore, an additional systematic uncertainty of $8\times 10^{-8}$ per ion per cycle ($3\times 10^{-7}$ total) was required to account for ion-ion interactions (based on off-line tests).
Figure 2: Fit of the overlapping peaks in the $A=153$ spectrum with a hyperEMG functional form.
According to ENSDF, 153Tm has a long-lived isomer at 43.2$\pm$0.2 keV [46] and 155Tm one at 41$\pm$6 keV [47]. 151Tm, 152Tm, and 154Tm respectively contain a $\sim 7$ s, a $5.2$ s, and a $3.3$ s isomer [48, 49, 50] for which the excitation energy was taken from NUBASE 2020 [51] to be 93$\pm$6 keV, -100$\pm$250 keV, and 70$\pm$50 keV. NUBASE 2020 also infers excitation energies for 149Tm (100#$\pm$50# keV) [52] and 150Tm (140#$\pm$140# keV) [53] from trends in neighbouring nuclei; the higher 671 keV, 5.2 ms isomer of 150Tm is not expected to be observable in this experiment due to its short half-life and no indication of it was observed. 157Lu also contains an unknown contribution from a 20.9$\pm$2.0 keV isomer (4.8 s) [54].
Since these isomeric contributions may shift the determined mass to appear heavier than for a peak containing only the ground state, the standard procedure outlined in Appendix B of AME2020 [55] was used to determine an appropriate correction to the mass and the corresponding increase in uncertainty. The ground state mass ($m_{gs}$) was obtained from the experimental mass ($m_{exp}$) and the previously reported excitation energy ($\Delta E$) through the relation $m_{\textrm{gs}}=m_{\textrm{exp}}-\Delta E/2$, and additional uncertainties of $0.290(\Delta E)$, and $\frac{1}{2}\sigma_{1}$ were added in quadrature with the experimental error, where $\sigma_{1}$ is the uncertainty in the reported excitation energy. The Mass Excesses and Ion Mass Ratio was thus corrected for all species likely containing an unresolved isomer (e.g. 13 keV for 153gTm and 12 keV for 155gTm).
The mass spectra of 149,150Tm strongly benefited from using mass selective re-trapping to improve the Tm-to-contaminant ratio by about three to four orders of magnitude depending on their separation, as shown in Figure 1 of reference [15]. Once applied, the peaks were observable above the background and could be fit.
@@
As seen in Figure 3, our measured mass values show overall agreement with the literature values from AME2020, including for 157Tm and 156Tm which are not known to have long-lived isomers or isobaric species that could impact our measurements. The two new mass values of 149,150Tm now close a gap in the long chain of previously measured neutron-deficient isotopes of Tm [18, 56, 57], with mass data now spanning from the proton emitter 147Tm at N=78 to stable 169Tm at N=100. Alongside the measured Tm isotopes, 149-154Er were also identified in the mass spectra. The masses of 150-154Er all lie within 1.5 $\sigma$ of the presently accepted values in the AME2020 and any known isomers are $>2.5$ MeV above the ground state and thus do not impact the measurements. After correcting for the isomer, our measurement of 157Lu yields a mass 56$\pm$29 keV heavier than the AME value.
@@
149gEr and 149mEr (bottom Figure 1) were fit using two hyperEMG functions calibrated to the shape of the 149Dy peak. The known $11/2^{-}$ isomeric state was delivered to TITAN at greater intensity and the ground state peak shows up as an excess in the left shoulder. Our 149gEr and 149mEr masses were found to be heavier by 2.9 $\sigma$ and 3.9 $\sigma$ respectively when compared to the literature value obtained via a measurement of 149mEr at the GSI ESR [14]. This results in a shift of the $S_{p}$ of 150Tm (see below).
Figure 3: Comparison of the measured Tm isotopes with the error bounds on the ground state given in the 2020 Atomic Mass Evaluation. The masses and error bars of measurements marked with a purple “x” or a blue triangle were corrected for unresolved long-lived isomers listed in ENSDF and NUBASE respectively.
@@
Table 1: List of isotopes, their respective calibrants (AME2020 values), measured ion mass ratios, and the mass excesses when referenced to the AME2020 values of the calibrant (uncertainty added in quadrature). The 150Tm mass from experiment #1 [15] is included at the bottom. The # indicates an extrapolated mass in the AME2020. Nuclides with a * have a long-lived isomer for which the excitation energy is reported in NUBASE and relies on alpha decay data, and nuclides with a ** rely on extrapolations. Nuclides with a contained an unresolved admixture of the ground state and an isomer in the peak for which the known excitation energy is listed in both ENSDF and NUBASE. The mass excess and the ion mass ratio of any peaks likely to contain an unresolved isomeric contribution were corrected by 0.5 $\times$ the listed excitation energy (listed in bold) and a corresponding additional uncertainty was incorporated (see Section III). The 155Tm peak has an unresolved 41(6)keV isomer [47], 153Tm has a 43.2$\pm$0.2 keV isomer [46], and 157Lu has a 20.9$\pm$2.0 keV isomer [54]. As described in the text, these values were corrected using the procedure outlined in the 2020 Atomic Mass Evaluation (AME2020) [55] to obtain a mass for the ground state and its uncertainty. The difference between the TITAN measurement and the AME2020 value is given with the AME and the measurement uncertainty added in quadrature.